# EE-LLM: Large-Scale Training and Inference of Early-Exit Large Language Models with 3D Parallelism

### 08 Dec 2023 v1

@@ 4.2.2 Further performance optimization
So far, we have analyzed training time and memory usage by model parameters. One remaining issue that we have not addressed is the activation memory overhead due to early-exit computation. Most notably, the early-exit logits for one microbatch have size $s\times b\times V$, where $s$ is the maximum sequence length, $b$ is the microbatch size, and $V$ is the vocabulary size. If the $i$-th stage has one early exit, then a vanilla implementation of training with early exits using the 1F1B pipeline schedule (as shown in Figure 3(b)) will cause a significant memory overhead of size $s\times b\times V\times(P-i+1)$, where $P-i+1$ is the number of in-flight microbatches [20] for Stage $i$.
@@
One can check that, with the above modification, our analysis in Section 4.2.1 about the training time overhead caused by early exits remains valid, with minor modifications. More specifically, the time overhead of one training iteration is still $k\times(f_{\mathsf{EE}}+b_{\mathsf{EE}})$; the only difference is that this whole overhead comes from the backward steps of the last microbatch (i.e. the third part of the critical path in Figure 3(c)). One can further reduce this overhead to $k\times b_{\mathsf{EE}}$ by moving the forward pass of the early-exit layer on each stage for each cool-down microbatch to the explicit bubble in front of the corresponding backward step (i.e. before communication with the next stage). With that said, our implementation does not include this modification, as it brings limited gains at the cost of complicating our codebase.
@@
Recall from Section 3.1 that, with EE-LLM, users can choose more expressive and powerful early-exit layers beyond the minimalistic structure. Similarly, more than one early exit can be added to each pipeline stage, which provides more flexible choices of exits during inference. These benefits, of course, come at the cost of higher overhead for training, and potentially for inference as well444 There is no clear answer to whether additional modules at early-exit layers will improve or hurt the overall inference speed. There is certainly higher overhead for the computation of each early-exit layer; on the other hand, higher flexibility and adaptivity of the early exits can potentially enable them to produce better outputs and get selected more often during inference, leading to overall faster generation of a complete sequence. For similar reasons, there is no clear positive or negative correlation between the number of early exits and the overall speed of generating a sequence.; with EE-LLM, users can conveniently choose the most suitable configurations for their own use cases.
@@
Figure 4: The forward time, backward time, and peak GPU memory of each pipeline stage for a standard 7B GPT Transformer, as well as its early-exit version (with one minimalistic early-exit layer added to each middle stage). Degrees of pipeline, tensor and data parallelism are 4, 1, and 1, respectively; the microbatch size is 2, global batch size is 128, and sequence length is 2048. Note that the forward computation of early-exit layers has been deferred to the backward steps, hence not included in “forward time” of the first plot, but in “backward time” of the second plot.
@@ 4.3 Advanced features
EE-LLM incorporates some advanced features that can potentially improve the training process, which are introduced below. We note that these are exploratory functionalities, and their practical benefits are still under testing.
4.3.1 Adaptive weights of early-exit losses
The weights of early-exit losses can be changing (rather than constant) during the training process, just like the learning rate or other hyperparameters. Setting the weights adaptively can offer a more fine-grained control of how gradients from multiple loss functions jointly impact the backbone and early/final output layers of the model. This idea has been adopted in some previous works [18], and we implement and offer such an option in EE-LLM as well.
@@ 4.3.2 Filling explicit bubbles with additional microbatches
To further leverage the explicit bubbles within the 1F1B pipeline schedule (i.e. the gray areas in Figure 5), we design and implement a novel approach of filling them with partial forward/backward computation of additional microbatches. This is primarily inspired by the idea from [30]: instead of designing a new, sophisticated pipeline schedule that has a lower bubble ratio, one may seek to fill the bubbles of an existing schedule with useful computation, which can lead to better resource utilization and faster training.
@@
One thing to note is that this approach changes the optimization semantics. From the perspective of stochastic optimization, we can prove formally that with such additional computation (and under certain conditions), one obtains an unbiased gradient estimate with reduced variance for the original training objective. We refer interested readers to Appendix A for more details about the methodology and analysis.
Figure 5: A visualization for our proposed method of filling pipeline bubbles with partial forward and backward computation of additional microbatches (i.e. the ones annotated as P1, P2, P3 and P4).
@@ 5.1 Main challenge: KV caching
Accelerating inference with early exiting is, in principle, orthogonal to (and thus compatible with) many other common techniques of acceleration [19], such as kernel fusion, FlashAttention [10], model quantization [50], among others. For autoregressive generation tasks, one exception is KV caching, i.e. saving keys and values of all attention layers for previously generated tokens, which accelerates sequence generation by reducing redundant computation (though at the cost of higher memory usage). This is obviously in contradiction with vanilla early-exit inference: if the current token is generated via early exiting, then its KV caches in later layers are missing, which hinders the generation of future tokens that go beyond the exiting layer of the current token.
This challenge has been well recognized in the literature, and several approaches have been recently proposed to resolve it. One approach [11, 36] is to copy the hidden states of the current token at the exiting layer to all later layers, which will be used to compute the keys and values at later attention layers. Despite its efficiency, this method obviously causes a deviation from the inference process that the model is trained to excel at, which can harm the output quality. Another solution [9] is to pre-specify the exiting layer for each future token, while ensuring that KV missing in previous tokens will not hinder generation of later tokens; with this approach, the ability of token-wise adaptive selection of exiting layers is inevitably lost. The third method [2, 43] is to store the hidden states of recent tokens that were generated with early exiting, and whenever KV missing happens, run a batch forward pass with the current and recent tokens to fulfill the KV caches. A variant of this approach, which we call KV recomputation, is implemented in EE-LLM; we defer its details to Section 6.3. Despite the batching effect of GPU computation, this approach still brings additional computational overhead, and might not even achieve any acceleration for inference on other hardware platforms, due to its high computational complexity.
@@ 5.2 A new solution: pipeline parallelism
We propose a novel solution that leverages a new type of pipeline parallelism during inference. The key idea is: in the process of inference with multiple pipeline stages, the following two processes run in parallel whenever the model decides to do early exiting for the current token at a certain early exit:
@@
See Figure 6 for a visualization of our proposed approach. Even though each token goes through a forward pass of the full model, the computation after the exiting layer is parallelized with the computation of later tokens, which is essentially how acceleration is achieved in this approach. An implementation of this method is provided in EE-LLM.
One potential limitation of the proposed method is that it requires multiple devices to facilitate pipeline parallelism (although parallelism within a single GPU or other device might be possible); moreover, the maximum possible speedup for generating each token is upper bounded by the number of pipeline stages, even if there is an early exit in the middle of the first stage.
Figure 6: A comparison between the standard full-model inference (top), and our pipeline-based approach of early-exit inference (bottom). Numbers in the blocks represent the tokens within one generated sequence.
@@ 6 Implementations
The implementation of EE-LLM is based on Megatron-LM [28], primarily extending Megatron-LM’s (1) model architectures, (2) pipeline scheduling, and (3) inference service to support the training and inference of early-exit LLMs. We introduce each of these aspects in more details below.
@@ 6.1 Model architectures
We have introduced a new class of models called EarlyExitGPTModel, which is the early-exit counterpart of GPTModel in the original model library of Megatron-LM. The model is constructed with a few other classes, including EarlyExitTransformerLayer, EarlyExitTransformer, and EarlyExitLanguageModel. EarlyExitTransformerLayer is a replacement for the original ParallelTransformerLayer in Megatron-LM. It adds an early-exit structure on top of the standard Transformer layer, which allows it to generate outputs for both the main network backbone and the early exit; for the latter, it returns a lazy loss function during training, or tokens during inference. This module supports various customizations of the early-exit structure; besides the minimalistic structure with an output embedding matrix and an (optional) output normalization layer, one might add e.g. a MLP or a complete Transformer layer. These additional structures can be combined in any desired manner and can be placed before or after the backbone part of this layer. On the other hand, EarlyExitTransformer and EarlyExitLanguageModel are mainly used to propagate the early-exit outputs to the top-level model. They are capable of stopping the forward computation at the early-exit layer and returning the intermediate outputs, which facilitates accelerated inference.
@@ 6.2 Pipeline scheduling
We have adjusted the existing 1F1B schedule for early-exit LLMs, as shown in Figure 3. To fill implicit bubbles and reduce GPU memory overhead, lazy loss functions of early-exit layers are returned together with outputs of the backbone network during forward steps. These lazy functions are not actually called until their corresponding auxiliary losses (see Section 4.1.1) are calculated in the backward steps. For the method of filling explicit bubbles proposed in Section 4.3.2, we have inserted partial forward/backward computation of additional microbatches into warm-up and cool-down phases of the 1F1B schedule. The number of inserted microbatches and partial forward/backward stages can be automatically calculated through the user-specified (estimate of) ratio between backward and forward time.
@@ 6.3 Inference service
For inference with pipeline parallelism (i.e. the pipeline-based approach proposed in Section 5.2), we have re-implemented the forward process. With our implementation, the first pipeline stage will wait for an exit signal from the early/final exits of all subsequent stages after its forward computation is completed. Each subsequent stage will send an exit signal and the output token to the first stage, if there is an exit within the stage that satisfies the exit condition. Upon receiving the signal and generated token, the first stage will immediately start the forward pass for generating the next token. With this implementation, regardless of the early-exit layers’ positions in subsequent stages, the inference service can immediately generate a token whenever early exiting happens on some stage, without waiting for the completion of the entire stage (except for the first stage).
For inference without pipeline parallelism, we have implemented a mechanism of KV recomputation (which is a variant of synchronized parallel decoding proposed recently in [2]). In this approach, we maintain a list of the most recent tokens that have missing KV caches in deep layers due to early exiting. During each forward pass, we include these early-exit tokens in the current forward pass, which allows for direct recomputation of the KV caches for these tokens and thus avoids the issue of missing KV caches. Acceleration of sequence generation is still achieved, thanks to the batching effects of GPU computation. To avoid the endless accumulation of early-exit tokens, we enforce a full-model forward pass whenever the number of early-exit tokens reaches a pre-specified value.
@@ 7 Experiments
This section provides an empirical evaluation of the training and inference efficiency achieved by EE-LLM. Our experiments for training are run on an 8-node cluster with 64 Nvidia A100-80GB GPUs, where each node has 8 GPUs. We use the pre-training data from Data-Juicer [5], which consists of 800 billion tokens. For inference, we use a server with 4 Nvidia A100-40GB GPUs.
@@ 7.1.1 Convergence of training losses
First of all, we conduct experiments to verify the convergence of training losses when training early-exit LLMs with EE-LLM. We first consider a 1.3B GPT Transformer, add one minimalistic early-exit layer (without layer normalization) to the 1/4 depth and the other to the 1/2 depth, set their early-exit loss weights to 1/4 and 1/2 respectively (while the final-exit loss has a weight of 1), and tie all input and output embedding matrices; a standard LLM of the same GPT architecture is also trained using the same hyperparameters and pre-training data. We further train a 7B early-exit model using similar configurations, except that early-exit loss weights are set to 0.1 and 0.2, and all embedding matrices are untied.
Figure 7 shows the convergence of early-exit and final-exit training losses (i.e. negative log-likelihood of next-token prediction) for both standard and early-exit LLMs. All loss curves decay at a similar pace, and unsurprisingly, the early-exit losses are slightly higher than the final-exit loss for each model. Interestingly, for 1.3B models, the final-exit losses of both standard and early-exit models largely overlap with each other, suggesting that optimizing for early-exit losses might not hurt the final output of the full model in our setting.
@@ 7.1.2 Training efficiency
In this experiment, we investigate the training overhead of early exits. Starting with a standard GPT Transformer of size ranging from 1.3B to 30B, we increase the number of early exits from 0 to 3. Minimalistic early exits are added one by one to specific locations in the following order: (1) to the 1/4 depth; (2) to the 1/2 depth; (3) to the hidden states right before the first Transformer layer (which is always located in the first pipeline stage). Based on the performance optimizations proposed in Section 4.2, when an early exit is inserted into the middle of two layers located on two consecutive pipeline stages, we always add it to the beginning of the latter stage. We set the global batch size to 2048, microbatch size to 2 (for 1.3B and 7B models) or 1 (for 13B and 30B models), and sequence length to 2048. The degree of data parallelism is set to 4 and fixed throughout, while tensor/sequence and pipeline parallelism degrees can take various values.

### 01 Feb 2024 v2

@@ 4.2.2 Further performance optimizations
So far, we have analyzed training time and memory usage by model parameters. Another major component of memory usage, namely the memory for gradients and optimizer states, can be bounded by the memory for model parameters multiplied by a universal constant. One remaining issue that we have not addressed is the activation memory overhead due to early-exit computation. Most notably, the early-exit logits for one microbatch have size $s\times b\times V$, where $s$ is the maximum sequence length, $b$ is the microbatch size, and $V$ is the vocabulary size. If the $i$-th stage has one early exit, then a vanilla implementation of training with early exits using the 1F1B pipeline schedule (as shown in Figure 3(b)) will cause a significant memory overhead of size $s\times b\times V\times(P-i+1)$, where $P-i+1$ is the number of in-flight microbatches [36] for Stage $i$.
@@
One can check that, with the above adjustment, our analysis in Section 4.2.1 about the training time overhead caused by early exits remains valid after minor modifications. More specifically, the time overhead of one training iteration is still $k\times(f_{\mathsf{EE}}+b_{\mathsf{EE}})$; the only difference is that this whole overhead comes from the backward steps of the last microbatch, i.e. the third part of the critical path in Figure 3(c). One can further reduce this overhead to $k\times b_{\mathsf{EE}}$ by moving the forward pass of the early-exit layer on each stage for each cool-down microbatch to the explicit bubble in front of the corresponding backward step (i.e. before communication with the next stage). With that said, our implementation does not include this modification, as it brings limited gains at the cost of complicating our codebase.
@@
Recall from Section 3.1 that, with EE-LLM, users can choose more expressive and powerful early-exit layers beyond the minimalistic structure. Similarly, more than one early exit can be added to each pipeline stage, which provides more flexible choices of exits during inference. These benefits, of course, come at the cost of higher overhead for training, and potentially for inference as well444 There is no clear answer to whether additional modules at early-exit layers will improve or hurt the overall inference speed. There is certainly higher overhead for the computation of each early-exit layer; on the other hand, higher flexibility and adaptivity of the early exits can potentially enable them to produce better outputs and get selected more often during inference, leading to overall faster generation of a complete sequence. For similar reasons, there is no clear positive or negative correlation between the number of early exits and the overall speed of generating a sequence.; with EE-LLM, users can conveniently choose the most suitable configurations for their own use cases. We refer interested readers to Appendix A for formal analysis of training efficiency in these general cases.
@@
Figure 4: The forward time, backward time, and peak GPU memory of each pipeline stage for a standard 7B GPT Transformer, as well as its early-exit version that has one minimalistic early-exit layer (without layer normalization) added to each middle stage. Degrees of pipeline, tensor and data parallelism are 4, 1, and 1, respectively; the microbatch size is 2, global batch size is 128, and sequence length is 2048. Note that the forward computation of early-exit layers has been deferred to the backward steps, hence not included in “forward time” of the first plot, but in “backward time” of the second plot.
@@ 4.3 Advanced features
EE-LLM incorporates some advanced features that can potentially improve the training process, which are introduced below. We note that these are exploratory functionalities, and formal investigation of their practical benefits is left for future work.
4.3.1 Non-constant weights of early-exit losses
The weights of early-exit losses in the training objective of Eq. (1) can be changing rather than constant during the training process, just like the learning rate or other hyperparameters. Allowing the weights to be changing can offer a more fine-grained control of how gradients from multiple loss functions jointly impact the backbone and early/final output layers of the model. This functionality has been implemented in EE-LLM.
One concrete option that we offer is warm-up. With this option, early-exit loss weights start at small values, and gradually increase with training iterations until reaching the pre-specified maximum values. This approach has been adopted in prior works [31]. The idea is to encourage the deep neural network to primarily optimize for the full-model output quality from the beginning of the training process, while the skill of early exiting is gradually acquired with minor or no negative impact on the final outputs of the full model.
Another option is cool-down, which does the opposite and decreases early-exit loss weights during the training process. This option is inspired by prior works [39, 69] that leverage early-exit losses for the purpose of regularizing the training process of deep neural networks. Such “deep supervision” provided by early-exit losses can stabilize and accelerate convergence of training, and potentially improve the intermediate features learned by the neural network. As early-exit loss weights gradually decay, the regularization gets weaker, so that the network becomes more focused on its primary objective, namely the final-output quality.
@@ 4.3.2 Filling explicit bubbles with additional microbatches
To further leverage the explicit bubbles within the 1F1B pipeline schedule, i.e. the gray areas in Figure 5, we design and implement a novel approach of filling them with partial forward/backward computation of additional microbatches. This is primarily inspired by the idea from [51]: instead of designing a new, sophisticated pipeline schedule that has a lower bubble ratio, one may seek to fill the bubbles of an existing schedule with useful computation, which can lead to better resource utilization and faster training.
@@
One thing to note is that this approach changes the optimization semantics. From the perspective of stochastic optimization, we can prove formally that with such additional computation and under certain conditions, one obtains an unbiased gradient estimate with reduced variance for the original training objective. We refer interested readers to Appendix B for more details about the methodology and analysis.
Figure 5: A visualization for our proposed method of filling pipeline bubbles with partial forward and backward computation of four additional microbatches, i.e. the ones annotated as P1, P2, P3 and P4.
@@ 5.1 Main challenge: KV caching
Accelerating inference with early exiting is, in principle, orthogonal to and compatible with many other common techniques of acceleration [32], such as kernel fusion, FlashAttention [15], model quantization [85, 80], among others. For autoregressive generation tasks, one exception is KV caching, i.e. saving keys and values of all attention layers for previously generated tokens, which accelerates sequence generation by reducing redundant computation (though at the cost of higher memory usage). This is obviously in contradiction with vanilla early-exit inference: if the current token is generated via early exiting, then its KV caches in later layers are missing, which hinders the generation of future tokens that go beyond the exiting layer of the current token.
This challenge has been well recognized in the literature, and several approaches have been recently proposed to resolve it. One approach [19, 42, 61] is to copy the hidden states of the current token at the exiting layer to all later layers, which will be used to compute the keys and values at later attention layers. Despite its efficiency, this method obviously causes a deviation from the inference process that the model is trained to excel at, which can harm the output quality. Another solution [13] is to pre-specify the exiting layer for each token, while ensuring that KV missing in previous tokens will not hinder generation of later tokens; with this approach, the ability of token-wise adaptive selection of exits is inevitably lost. The third method [3, 70] is to store the hidden states of recent tokens that were generated with early exiting, and whenever KV missing happens, run a batch forward pass with the current and recent tokens to fulfill the KV caches. A variant of this approach, which we call KV recomputation, is implemented in EE-LLM; we defer its details to Section 6.3. Despite the batching effect of GPU computation, this approach still brings additional computational overhead, and might not even achieve any acceleration for inference on other hardware platforms, due to its high computational complexity.
@@ 5.2 A new solution: pipeline parallelism
We propose a novel solution that leverages a new type of pipeline parallelism during inference. The key idea is that, in the process of inference with multiple pipeline stages, the following two processes run in parallel whenever the model decides to do early exiting for the current token at a certain early exit:
@@
See Figure 6 for a visualization of our proposed approach. Even though each token essentially goes through a forward pass of the full model, the computation after the exiting layer is parallelized with the computation of later tokens, which is how acceleration is achieved in this approach. It can be checked that the inference latency for generating one token at a certain exit matches exactly the time needed for the forward computation before returning an output at that exit, unless the selected exit is located in the middle of the first pipeline stage, in which case generation of the next token has to wait until the forward pass of the first stage for the current token is completed. Note that this is true not just in practice but also for the theoretical time complexity, without relying on the batching effect of GPU computation like KV recomputation does.
An implementation of this pipeline-based inference method is provided in EE-LLM. One potential limitation of the proposed method is that it requires multiple devices to facilitate pipeline parallelism, although parallelism within a single GPU or other device might be possible with more advanced implementation.
Figure 6: A comparison between the standard full-model inference (top), and our pipeline-based approach of early-exit inference (bottom). Numbers in the blocks represent the tokens within one generated sequence. Here, we consider the special case where each early exit is located at the end of a certain pipeline stage, just for simplicity of visualization.
@@ 6 Implementations
The implementation of EE-LLM is based on Megatron-LM [49], primarily extending Megatron-LM’s model architectures, pipeline scheduling, and inference service to support the training and inference of early-exit LLMs. We introduce each of these aspects in more details below.
@@ 6.1 Model architectures
We have introduced a new class of models called EarlyExitGPTModel, which is the early-exit counterpart of GPTModel in the original model library of Megatron-LM. The model is constructed with a few other classes, including EarlyExitTransformerLayer, EarlyExitTransformer, and EarlyExitLanguageModel. EarlyExitTransformerLayer is a replacement for the original ParallelTransformerLayer in Megatron-LM. It adds an early-exit structure on top of the standard Transformer layer, which allows it to generate outputs for both the main network backbone and the early exit; for the latter, it returns a lazy loss function during training, or tokens during inference. This module supports various customizations of the early-exit structure; besides the minimalistic structure with an output embedding matrix and an optional output normalization layer, one might add e.g. a MLP or a complete Transformer layer. These additional structures can be combined in any desired manner and can be placed before or after the backbone part of this layer. On the other hand, EarlyExitTransformer and EarlyExitLanguageModel are mainly used to propagate the early-exit outputs to the top-level model. They are capable of stopping the forward computation at the early-exit layer and returning the intermediate outputs, which facilitates accelerated inference.
@@ 6.2 Pipeline scheduling
We have adjusted the existing 1F1B schedule for early-exit LLMs, as shown in Figure 3. To fill implicit bubbles and reduce GPU memory overhead, lazy loss functions of early-exit layers are returned together with outputs of the backbone network during forward steps. These lazy functions are not actually called until their corresponding auxiliary losses (cf. Section 4.1.1) are calculated in the backward steps. For the method of filling explicit bubbles proposed in Section 4.3.2, we have inserted partial forward/backward computation of additional microbatches into warm-up and cool-down phases of the 1F1B schedule. The number of inserted microbatches and partial forward/backward stages can be automatically calculated through the user-specified (estimate of) ratio between backward and forward time.
@@ 6.3 Inference service
For inference with pipeline parallelism, i.e. the pipeline-based approach proposed in Section 5.2, we have re-implemented the forward process. With our implementation, the first pipeline stage will wait for an exit signal from the early/final exits of all subsequent stages after its forward computation is completed. Each subsequent stage will send an exit signal and the output token to the first stage, if there is an exit within the stage that satisfies the exit condition. Upon receiving the signal and generated token, the first stage will immediately start the forward pass for generating the next token. With this implementation, regardless of the early-exit layers’ positions in subsequent stages, the inference service can immediately generate a token whenever early exiting happens on some stage, without waiting for the completion of the entire stage (except for the first stage).
For inference without pipeline parallelism, we have implemented a mechanism of KV recomputation, which is a variant of synchronized parallel decoding proposed recently in [3]. In this approach, we maintain a list of the most recent tokens that have missing KV caches in deep layers due to early exiting. During each forward pass, we include these early-exit tokens in the current forward pass, which allows for direct recomputation of the KV caches for these tokens and thus avoids the issue of missing KV caches. Acceleration of sequence generation is still achieved, thanks to the batching effects of GPU computation. To avoid the endless accumulation of early-exit tokens, we enforce a full-model forward pass whenever the number of early-exit tokens reaches a pre-specified value.
@@ 7 Experiments
This section provides an empirical evaluation of the training and inference efficiency achieved by EE-LLM.
@@ 7.1 Training
In the following experiments, we empirically investigate the convergence of training early-exit models with EE-LLM, as well as the training efficiency of EE-LLM for early-exit LLMs up to an unprecedented scale of 30B. This scale is only limited by the hardware resources available to us, namely an 8-node cluster with 8 Nvidia A100-80GB GPUs in each node and hence 64 GPUs in total. We use a subset of the pre-training data [^5] https://github.com/alibaba/data-juicer/blob/main/configs/data_juicer_recipes/README.md provided by Data-Juicer [9], which consists of 800 billion tokens.
@@ 7.1.1 Convergence of training losses
First of all, we conduct experiments to verify the convergence of training losses when training early-exit LLMs with EE-LLM. We first consider a 1.3B GPT Transformer with 24 layers, add one minimalistic early-exit layer without layer normalization to the 1/4 depth and the other to the 1/2 depth, set their early-exit loss weights to 1/4 and 1/2 respectively (while the final-exit loss has a weight of 1), and tie all input and output embedding matrices. A standard LLM of the same GPT architecture is also trained using the same hyperparameters and pre-training data. We further train a 7B early-exit model with 32 layers using similar configurations, except that early-exit loss weights are set to 0.1 and 0.2, and all embedding matrices are untied. A 7B standard model is trained similarly. In all cases, we use the Adam optimizer [35] with hyperparameters $\beta_{1}=0.9,\beta_{2}=0.95,\epsilon=10^{-8}$, and a cosine schedule for the learning rate, with a maximum value of $3\times 10^{-4}$. The batch size and sequence length are both set to 2048.
Figure 7 shows the convergence of early-exit and final-exit training losses, i.e. negative log-likelihood of next-token prediction, for both standard and early-exit LLMs. All loss curves decay at a similar pace, and unsurprisingly, the early-exit losses are slightly higher than the final-exit loss for each model. Interestingly, the final-exit loss curve of each early-exit model is close to (or even slightly below) that of the standard model, suggesting that optimizing for early-exit losses might not hurt the final output of the full model in our setting. We also observe more spikes in the loss curves for the 7B model than for the 1.3B models, possibly due to the following reasons: (1) we choose to untie the embedding matrices and use smaller early-exit loss weights, both of which incur weaker regularization for the training process; (2) layer normalization, which is known to stabilize training of LLMs, is not included in the minimalistic early-exit layers of our 7B model.
@@ 7.1.2 Training efficiency
In this experiment, we investigate the training overhead of early exits. Starting with a standard GPT Transformer of size ranging from 1.3B to 30B, we increase the number of early exits from 0 to 3. Minimalistic early exits are added one by one to specific locations in the following order: (1) to the 1/4 depth; (2) to the 1/2 depth; (3) to the hidden states right before the first Transformer layer, which is always located in the first pipeline stage. Based on the performance optimizations proposed in Section 4.2, when an early exit is inserted into the middle of two layers located on two consecutive pipeline stages, we always add it to the beginning of the latter stage. We set the global batch size to 2048, microbatch size to 2 (for 1.3B and 7B models) or 1 (for 13B and 30B models), and sequence length to 2048. The degree of data parallelism is set to 4 and fixed throughout, while tensor/sequence and pipeline parallelism degrees can take various values.